OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units long. To the nearest whole number, what is the length of the hypotenuse? 8 units 9 units 10 units 11 units

OpenStudy (anonymous):

8

OpenStudy (angela210793):

h^2=4^2+7^2=16+49=65 h=sqr65~8

OpenStudy (lalaly):

by Pythagaras Theorem hypotenuse = square root of (4^2 + 7^2) = 8 units

OpenStudy (aravindg):

8

OpenStudy (anonymous):

hypotenuse^2=(4)^2+(7)^2 =16+49 =65 hypotenuse=8

OpenStudy (aravindg):

ya

OpenStudy (anonymous):

ya 8